OpenStudy (pokemon23):

The larger of two numbers is 1 more than 3 times the smaller. The difference between 8 times the smaller an 2 times the larger is 10. Find the numbers. The larger of two numbers is 1 more than 3 times the smaller. The difference between 8 times the smaller an 2 times the larger is 10. Find the numbers. @Mathematics

OpenStudy (pokemon23):

Help me

OpenStudy (heisenberg):

can you set up the equations?

OpenStudy (pokemon23):

you back

OpenStudy (pokemon23):

let me try

OpenStudy (pokemon23):

idk get it

OpenStudy (anonymous):

The numbers are 6 and 19.

OpenStudy (pokemon23):

what the equation?

OpenStudy (anonymous):

x=3y+1, 8y-2x+10 Solve them by elimination method.

OpenStudy (anonymous):

Sorry! It's 8y-2x=10

OpenStudy (pokemon23):

idk get it

OpenStudy (pokemon23):

ty guys

OpenStudy (aravindg):

8y-2x=10

OpenStudy (pokemon23):

how do I simplify the equation

OpenStudy (aravindg):

u see x=1+3y

OpenStudy (pokemon23):

yes what about it

OpenStudy (aravindg):

and 8y-2x=10

OpenStudy (aravindg):

solve the two eans to get x and y

OpenStudy (aravindg):

note x is the larger no. and y the smaller

OpenStudy (aravindg):

u will get 6 and 19

OpenStudy (aravindg):

hapy?

OpenStudy (pokemon23):

nope? I still don't get it?

OpenStudy (aravindg):

??

OpenStudy (aravindg):

u dont know to solve eqn in 2 variables??

OpenStudy (pokemon23):

so the whole equation is 8y-2x=10

OpenStudy (aravindg):

no

OpenStudy (aravindg):

we have 2 eqns

OpenStudy (aravindg):

x-3y=1

OpenStudy (aravindg):

8y-2x=10

OpenStudy (aravindg):

multiply eqn 1 by 2

OpenStudy (aravindg):

2x-6y=2

OpenStudy (aravindg):

add to 2nd eqn

OpenStudy (aravindg):

we get 2y=12

OpenStudy (aravindg):

therefore y=6

OpenStudy (aravindg):

understood?

OpenStudy (pokemon23):

i guess i'm confuse because i can't see the steps